# Self-organizing Multiagent Target Enclosing under Limited Information and Safety Guarantees

### v1

@@ IV Simulation Results
In this section, we demonstrate the efficacy of the proposed control law in autonomously organizing the pursuers to enclose a stationary target. In the following results, the target is stationary at the position $[x_{T}\;y_{T}]=[0,0]^{\top}$ m. All pursuers start at a different distances from the target and move at a constant speed of $40$ m/s and. The initial position of pursuers are denoted by square markers in the subsequent trajectory plots. The desired proximity from the target is set as $r_{d}=100$ m for all the simulation results. The controller gains and other parameters are selected as follows: $K_{i}=10$, $\lambda_{i}=0.9$, $\eta_{i}=70000$, $\Delta=100$, and $r_{s}=50$ m.
@@
Figure 10: Collision-free self-organising target enclosing with 6 pursuers.
The first set of results illustrates the scenario in which six pursuers start away from the target, as depicted in Figure 10. The initial arrangement of the pursuers around the target in a clockwise direction is given as $P_{2}-P_{1}-P_{3}-P_{4}-P_{5}-P_{6}$, as shown by square markers in the figure. It is observed that the final arrangement of the pursuers on the desired proximity (as depicted by circular markers) is given as, $P_{2}-P_{3}-P_{1}-P_{4}-P_{6}-P_{5}$. This indicates that only $P_{3}$ and $P_{6}$ encounter a nearest colliding pursuer and reorganize to avoid collision. Meanwhile, the $P_{1},P_{2},P_{4},andP_{5}$ move directly to the desired proximity without encountering the nearest colliding pursuer. The profiles of the error variables and sliding manifold are depicted Figure 9(b), where the sliding manifold converges to zero within a finite time. The profiles of $S_{i}$ for only $P_{3}$ and $P_{6}$ depict sudden deviations at different instances of time, which represent the instance of activation/deactivation of the collision avoidance component of the control input. Note that whenever collision avoidance is activated, the range error diverges from zero momentarily to ensure safety.
Figure 9(c) depicts the profiles of lateral accelerations of the pursuers converging to the same constant value at a steady state, as all the pursuers eventually converge to the same loiter circle of radius of desired proximity. The lateral acceleration profiles of $P_{3}$ and $P_{6}$ vary from the other pursuers due to the activation of the collision avoidance component. Finally, the profiles of inter-pursuer variables are shown in Figure 9(d), where the profiles of $\psi_{ij}$ change sign crossing zero for only $P_{3}$ and $P_{6}$. This shows that these pursuers have autonomously reorganized from their initial configurations, where the profiles $r_{ij}^{\min}$ depict that the distance between the two closest pursuers never reaches zero. Further, at steady state, all the pursuers have $r_{ij}^{\min}>r_{s}$, which implies that all the pursuers only converge to the desired proximity when the sensing range of no two pursuers intersect. Therefore, all the pursuers have only the attractive component of the control input activated at the steady state.
Figure 11 illustrates the second set of results where initially, four pursuers ($P_{1}$ to $P_{4}$) are already moving on the circle of desired proximity when two other pursuers ($P_{5}$ and $P_{6}$) start away from the desired proximity. From the Figures 10(a), 10(b) and 10(c), it is evident that $P_{5}$ reaches the desired proximity without encountering a nearest colliding pursuer and autonomously organizing on the free space on the enclosing shape. While $P_{5}$ itself becomes the nearest colliding pursuer for $P_{6}$, resulting in $P_{6}$ taking a larger route to converge to the desired proximity behind $P_{6}$. This is verified by Figure 10(d), where only the profiles related to $P_{6}$ exhibit abrupt variation compared to other pursuers, indicating to activation of the repulsive component of the control input. Therefore, $P_{5}$ and $P_{6}$ are able to autonomously join the enclosing formation without colliding with other pursuers.
Furthermore, we offer video demonstrations of our simulation results, accessible at this link. The videos titled “results_1.mp4” and “results_2.mp4” showcase the simulation outcomes corresponding to the scenarios presented in the paper. Additionally,“results_3.mp4” and “results_4.mp4” illustrate the self-organizing target enclosing in large swarms with 10 and 20 pursuers, respectively.
@@ V Conclusions
In this paper, we proposed guidance laws for multiple pursuers to safely and autonomously organize at the desired proximity to encircle a stationary target, using only the relative information. In our approach, the agents are free to converge to any point on the enclosing shape as long as they maintain a safe distance between them, which may result in the pursuers not requiring a pre-determined formation structure and autonomously organizing on the enclosing shape. Our guidance strategy ensures that each pursuer selects at most one other pursuer based on certain collision conditions. We show that preventing collision with the selected other pursuer is sufficient to avoid collisions with all the other pursuers. To tailor the desired switching behavior required, we proposed a potential function such that the attractive component of the potential is activated all the time. Meanwhile, the repulsive component is activated whenever the pursuer encounters a nearest colliding pursuer. Further, we utilized sliding mode control to develop the reaching law, ensuring finite time convergence of the sliding manifold, followed by asymptotic convergence of the error to zero. The resulting guidance law enables pursuers to move towards the desired proximity along collision-free paths, offering increased maneuvering flexibility, requiring minimal information, and demonstrating robustness to the motion of other pursuers. Validation of the guidance laws with different numbers of pursuers attests to the merits of the proposed guidance algorithm. Potential future extensions of this work may involve incorporating a moving target, exploring three-dimensional spaces, and integrating obstacle avoidance techniques.

### v2

@@ IV Simulation Results
In this section, we demonstrate the efficacy of the proposed control law in autonomously organizing the pursuers to enclose a stationary target. In the following results, the target is stationary at the position $[x_{T}\;y_{T}]=[0,0]^{\top}$ m. All pursuers start at different distances from the target and move at a constant speed of $40$ m/s. The initial positions of pursuers are denoted by square markers in the subsequent trajectory plots. The desired proximity from the target is set as $r_{d}=100$ m for all the simulation results. The controller gains and other parameters are selected as follows: $K_{i}=10$, $\lambda_{i}=0.9$, $\eta_{i}=70000$, $\Delta=100$, and $r_{s}=50$ m.
@@
Figure 7: Collision-free self-organizing target enclosing with 6 pursuers.
@@
The first set of results illustrates the scenario in which six pursuers start away from the target, as in Figure 7. The initial arrangement of the pursuers around the target in a clockwise direction is given as $P_{2}-P_{1}-P_{3}-P_{4}-P_{5}-P_{6}$, as shown by square markers in the figure. It is observed that the final arrangement of the pursuers on the desired proximity (as depicted by circular markers) is given as, $P_{2}-P_{3}-P_{1}-P_{4}-P_{6}-P_{5}$. This indicates that only $P_{3}$ and $P_{6}$ encounter a nearest colliding pursuer and reorganize to avoid collision. Meanwhile, the $P_{1},P_{2},P_{4},$ and $P_{5}$ move directly to the desired proximity without encountering the nearest colliding pursuer. The profiles of the error variables and sliding manifold are depicted Figure 7(b), where the sliding manifold converges to zero within a finite time. The profiles of $S_{i}$ for only $P_{3}$ and $P_{6}$ depict sudden deviations at different instances of time, which represent instances of activation/deactivation of the collision avoidance component in the control input. Note that whenever collision avoidance is activated, the range error diverges from zero momentarily to ensure safety.
Figure 7(c) depicts the profiles of lateral accelerations of the pursuers converging to the same constant value at a steady state, as all the pursuers eventually converge to the same loiter circle of radius of desired proximity. The lateral acceleration profiles of $P_{3}$ and $P_{6}$ vary from the other pursuers due to the activation of the collision avoidance component. Finally, the profiles of inter-pursuer variables are shown in Figure 7(d), where the profiles of $\psi_{ij}$ change sign crossing zero for only $P_{3}$ and $P_{6}$. This shows that these pursuers have autonomously reorganized from their initial configurations, where the profiles $r_{ij}^{\min}$ depict that the distance between the two closest pursuers never reaches zero. Further, it can be observed that all the pursuers ensure $r_{ij}^{\min}>{\epsilon_{i}}^{*}\;\forall t\geq 0$ (as depicted by the black dotted line on $r_{ij}$ profiles in Figure 7(d)) to maintain a safe distance with the pursuers converging to the desired proximity with only the attractive component the potential function in the control input, as presented in the controller design.
Figure 8 illustrates the second set of results where initially, four pursuers ($P_{1}$ to $P_{4}$) are already moving on the circle of desired proximity when two other pursuers ($P_{5}$ and $P_{6}$) start away from the desired proximity. From the Figures 8(a), 8(b) and 8(c), it is evident that $P_{5}$ reaches the desired proximity without encountering a nearest colliding pursuer and autonomously organizing on the free space on the enclosing shape. While $P_{5}$ itself becomes the nearest colliding pursuer for $P_{6}$, resulting in $P_{6}$ taking a larger route to converge to the desired proximity behind $P_{6}$. This is verified by Figure 8(d), where only the profiles related to $P_{6}$ exhibit abrupt variation compared to other pursuers, indicating the activation of the repulsive component of the control input. Therefore, $P_{5}$ and $P_{6}$ are able to autonomously join the enclosing formation without colliding with other pursuers. Furthermore, we offer video demonstrations of our simulation results, accessible at https://t.ly/x533O. The videos titled “results_1.mp4” and “results_2.mp4” showcase the simulation outcomes corresponding to the scenarios presented in the paper. Additionally,“results_3.mp4” and “results_4.mp4” illustrate the self-organizing target enclosing in large swarms with 10 and 20 pursuers, respectively.
@@ V Conclusions
In this paper, we proposed guidance laws for multiple pursuers to safely and autonomously organize at the desired proximity to encircle a stationary target, using only the relative information. In our approach, the agents are free to converge to any point on the enclosing shape as long as they maintain a safe distance between them, which may result in the pursuers not requiring a pre-determined formation structure and autonomously organizing on the enclosing shape. Our guidance strategy ensures that each pursuer selects at most one other pursuer based on certain collision conditions. We show that preventing collision with the selected other pursuer is sufficient to avoid collisions with all the other pursuers. To tailor the desired switching behavior required, we proposed a potential function such that the attractive component of the potential is activated all the time. Meanwhile, the repulsive component is activated whenever the pursuer encounters a nearest colliding pursuer. Further, we utilized sliding mode control to develop the reaching law, ensuring finite time convergence of the sliding manifold, followed by asymptotic convergence of the error to zero. The resulting guidance law enables pursuers to move towards the desired proximity along collision-free paths, offering increased maneuvering flexibility, requiring minimal information, and demonstrating robustness to the motion of other pursuers. Validation of the guidance laws with different numbers of pursuers attests to the merits of the proposed guidance algorithm. Incorporating a moving target and consideration of actual vehicle models, possibly in 3D space, could be interesting to pursue in the future. Additionally, the current work can be extended to include an in-depth investigation of the switching behavior and the effect of dynamic obstacles in the environment.